Grade 12Mechanics

a particle is projected at an angle 30 degree with the horizontal with speed 20m/s.how high will it strike a wall 8 under root 3m away from point of projection?

nikki yadav
8 Years agoGrade 12

2 Answers

Nitesh
8 Years ago
On the basis of teacher Trajectory equationY=xtan(theta)-gx²/2u²cos²(theta)So Y=(3*1/√3) - (10*9/2*300)Y=1.58m
Subha
6 Years ago
Y=tan(theta) x  --  g^2x^2/2u^2cos^2theta
Y=tan30°×8√3  --  10× (8√3) ^2/2×20^2×cos^2 30°
Y=8-3.2
Y=4.8m
This is the correct answer